heart93:

Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d

jhonyy9:

f(x) = ?

lilianamendez2:

it say f(x)=x+a/b

Destiny564:

a=d=4b=c=2f(x)=x+4/2 g(x)=2x-4